If the length of a rectangle is 3 yards more than twice its width, and the perimeter of the rectangle is 36 yards, find the dimensions of the rectangle
Answer by mananth(16949)

let width be x
length = 2x+3
..
Perimeter = 2(L+W)
2*(x+2x+3)=36
2(3x+3)=36
6x+6=36
6x=30
x= 5 yards the width
2x+3 = 10+3 = 13 yards the length